# Planar Juggling of a Devil-Stick using Discrete VHCs

### v1

@@ Abstract
Index Terms:
Devil-stick, impulse controlled Poincaré map, nonprehensile manipulation, orbital stabilization, virtual holonomic constraints.
@@ I Introduction
We consider the problem of planar juggling of a devil-stick, which represents a nonprehensile manipulation task using intermittent impulsive inputs. The dynamics and control of juggling point objects, described by position alone, has been considered in [1, 2, 3, 4, 5, 6], with many works modeling the combined object-robot system with associated unilateral constraints and impact laws. In contrast to point objects, juggling a devil-stick [7, 8, 9, 10] requires modeling orientation in addition to position.
The dynamics and closed-loop control design for planar symmetric juggling appeared in [8, 9]. A coordinate transformation was used to transform the orbital stabilization problem to one of fixed-point stabilization using a Poincaré map in a rotating reference frame. Consequently, the dynamic model only permitted orbits with states that are symmetric about the vertical axis.
In this paper, we use the notion of discrete virtual holonomic constraints (DVHCs), introduced in [11], to define the trajectory of the devil-stick at discrete instants directly in the inertial frame. Similar to VHCs [12, 13, 14, 15], DVHCs are geometric constraints on the generalized coordinates that avoid the need for tracking a time-varying reference trajectory. However, while VHCs are enforced using continuous feedback such that they are satisfied for all time, DVHCs are enforced using discrete feedback and are satisfied only at the instants when impulsive inputs are applied. The work in this paper extends the concept of DVHCs from rotations to oscillations of the passive coordinate for the dynamical system in [11]. Just like the zero dynamics induced by VHCs, the discrete zero dynamics (DZD) induced by DVHCs can be used to infer system stability. The DZD derived in this paper demonstrates that the dynamics of planar devil-stick juggling is far richer than that presented in [8].
The main contributions of this work are highlighted below, particularly in comparison to [8]:
@@
  • It is not assumed a priori that the devil-stick is juggled between configurations that are symmetric in both position and velocity about the vertical axis.
  • The system permits infinitely many 2-periodic orbits, i.e., juggling motions, which are not necessarily symmetric about the vertical axis.
  • The symmetric juggling motion in [8] is shown to be a special case of the family of orbits generated by the DVHCs chosen in this paper.
The impulse controlled Poincaré map (ICPM) approach [16], which was used to stabilize the fixed point associated with planar symmetric juggling in the rotating reference frame in [9] is adapted to stabilize any 2-periodic juggling motion.
This paper is organized as follows. In Section II, we define the planar juggling problem and present the hybrid dynamics of the system. In Section III, we present our choice of DVHCs and a control design that enforces them. In Section IV, we derive the DZD, which provides conditions for stable, periodic juggling. In Section V, we describe a method to stabilize a desired juggling motion. In Section VI, we validate the approach using simulations.
@@ II-A System Description
Consider the three-DOF devil-stick in the vertical $xy$ plane in Fig.1. Its configuration is described by the independent generalized coordinates $(h_{x},h_{y},\theta)$. It is juggled using the control inputs $I$ and $r$, where $I$ is the impulse applied normal to the devil-stick and $r$ is the distance of its point of application from the center-of-mass. The devil-stick is juggled between the orientations $\theta^{\mathrm{odd}}\in(0,\pi/2)$ and $\theta^{\mathrm{even}}\in(\pi/2,\pi)$, which are chosen a priori. Therefore, $\theta\in[\theta^{\mathrm{odd}},\theta^{\mathrm{even}}]$ $\forall t$ during the juggling task. Impulsive inputs are applied alternately at $\theta=\theta^{\mathrm{odd}}$ and $\theta=\theta^{\mathrm{even}}$, and the devil-stick falls freely under gravity between successive applications of impulsive inputs, making the resulting system dynamics hybrid. Since there is one fewer control input than the number of generalized coordinates, the system is underactuated with one passive DOF. The vector of generalized coordinates is
@@
where $v\in\mathbb{R}^{2}$ and $\omega\in\mathbb{R}$.
@@ II-B Hybrid Dynamic Model
The hybrid dynamic model captures a single impulsive actuation [^1] Impulsive inputs cause no change in position coordinates and finite jumps in velocities [16]., followed by a flight phase under gravity. Using $k$ to represent instants immediately prior to applications of impulsive inputs, the dynamics of the active positions and velocities are [11]
@@
The hybrid dynamics of the passive coordinate and velocity are given by [11]
@@ III-A Constraints on Positions
We use discrete virtual holonomic constraints (DVHCs) [11] to define the trajectory of the center-of-mass of the devil-stick in terms of its orientation at the instants $k$ that impulsive inputs are applied on the devil-stick:
@@
where $\Phi:S^{1}\rightarrow\mathbb{R}^{2}$ is given by
$\Phi(\theta_{k})=\begin{bmatrix}\alpha\tan\theta_{k}\\ \beta\end{bmatrix}$ (12)
where $\alpha,\beta\in\mathbb{R}^{+}$ are parameters that can be chosen arbitrarily. It follows from (5) that the above DVHC places the center-of-mass of the devil-stick at $h(k)=\begin{bmatrix}\alpha\tan\theta^{\mathrm{odd}}&\beta\end{bmatrix}^{T}$ when $k$ is odd, and $h(k)=\begin{bmatrix}\alpha\tan\theta^{\mathrm{even}}&\beta\end{bmatrix}^{T}$ when $k$ is even. From the choices of $\theta^{\mathrm{odd}}$ and $\theta^{\mathrm{even}}$ in Section II-A, it follows that $h_{x}(k)>0$ when $k$ is odd, and $h_{x}(k)<0$ when $k$ is even. In particular, if $\theta^{\mathrm{even}}=\pi-\theta^{\mathrm{odd}}$, i.e., impulsive inputs are applied when the orientation of the devil-stick is symmetric about the vertical $y$ axis, $h_{x}(k+1)=-h_{x}(k)$ $\forall k$. The value of $h_{y}(k)$ is constant $\forall k$.
Remark 1
The results for planar symmetric juggling in [8] are a special case for the DVHC in (12) - see Section IV-C. Using different DVHCs, different juggling motions may be obtained; e.g., the value of $h_{y}(k)$ can be varied with $k$ by replacing $\beta$ in (12) by some function of $\theta_{k}$.
@@ III-B Constraints on Velocities
Using (12) in the above equation and simplifying,
$\begin{split}\Psi(\theta_{k},\omega_{k})=\frac{(-1)^{k}\omega_{k}}{{\Delta\theta^{*}}}&\begin{bmatrix}\alpha(\tan\theta_{k}-\tan\theta_{k+1})\\ 0\end{bmatrix}\\ -&\begin{bmatrix}0\\ \dfrac{(-1)^{k}g{\Delta\theta^{*}}}{2\omega_{k}}\end{bmatrix}\end{split}$ (21)
@@ III-C Control Design Enforcing the DVHC
We seek the control inputs that enforce the DVHC. We first treat $\delta_{k}$ as an input for simplicity. Using (11) and (19) in (3), we obtain
@@
We define $\eta_{k}\triangleq\Phi(\theta_{k+1})-\Phi(\theta_{k})$ to rewrite (22) as
$\begin{split}(\lambda-1)\rho_{k}+\eta_{k}=&[D\rho_{k}+\Psi(\theta_{k},\omega_{k})]\delta_{k}\\ &+\begin{bmatrix}-\sin\theta_{k}\\ \cos\theta_{k}\end{bmatrix}\frac{I_{k}\delta_{k}}{m}+\begin{bmatrix}0\\ -\frac{1}{2}g\delta_{k}^{2}\end{bmatrix}\end{split}$ (24)
whose $x$ and $y$ components can be expressed separately as
@@
where $\rho_{k}=\begin{bmatrix}\rho_{xk}&\rho_{yk}\end{bmatrix}^{T}$, $D\rho_{k}=\begin{bmatrix}D\rho_{xk}&D\rho_{yk}\end{bmatrix}^{T}$, $\eta_{k}=\begin{bmatrix}\eta_{xk}&\eta_{yk}\end{bmatrix}^{T}$ and $\Psi=\begin{bmatrix}\Psi_{x}&\Psi_{y}\end{bmatrix}^{T}$. Since $\sin\theta_{k}\neq 0$ for any $\theta_{k}$ in (5), we eliminate $I_{k}$ between (25a) and (25b) to obtain
@@
The time-of-flight $\delta_{k}$ is obtained by solving the above equation. The value of $I_{k}$ is then obtained from (25a) as
@@
It can be shown that [11]
$D\rho_{k+1}=\frac{(\lambda-1)\rho_{k}}{\delta_{k}}$ (29)
@@
$\delta_{k}$ $=\frac{(-1)^{k}2\omega_{k}\alpha}{g{\Delta\theta^{*}}}\left(1-\frac{\tan\theta_{k+1}}{\tan\theta_{k}}\right)$ (30)
$I_{k}$ $=\frac{(-1)^{k}m}{\cos\theta_{k}}\left[\frac{\omega_{k}\alpha}{{\Delta\theta^{*}}}\left(1-\frac{\tan\theta_{k+1}}{\tan\theta_{k}}\right)+\frac{g{\Delta\theta^{*}}}{2\omega_{k}}\right]$ (31)
$r_{k}$ $=\frac{(-1)^{k+1}J{\Delta\theta^{*}}\cos\theta_{k}}{m\alpha\left(1-\dfrac{\tan\theta_{k+1}}{\tan\theta_{k}}\right)}$ (32)
@@ IV-A Derivation
The discrete zero dynamics (DZD) provides the dynamics of passive coordinate $\theta$ and passive velocity $\omega$ when the DVHCs are satisfied, and system trajectories evolve such that $\rho_{k}\equiv 0\Rightarrow h(k)\equiv\Phi(\theta_{k})$ and $D\rho_{k}\equiv 0\Rightarrow v(k)\equiv\Psi(\theta_{k},\omega_{k})$. It follows from (19) and (21) that
$v_{x}(k)$ $=\frac{(-1)^{k}\omega_{k}\alpha}{{\Delta\theta^{*}}}(\tan\theta_{k}-\tan\theta_{k+1})$ (33a)
$v_{y}(k)$ $=-\dfrac{(-1)^{k}g{\Delta\theta^{*}}}{2\omega_{k}}$ (33b)
Similarly, it follows from (7), (19) and (21) that
$v_{x}(k+1)$ $=\frac{(-1)^{k}\omega_{k+1}\alpha}{{\Delta\theta^{*}}}(\tan\theta_{k}-\tan\theta_{k+1})$ (34a)
$v_{y}(k+1)$ $=\dfrac{(-1)^{k}g{\Delta\theta^{*}}}{2\omega_{k+1}}$ (34b)
The input $I_{k}$ can be eliminated between the two equations in (4) to obtain
@@
$\frac{g{\Delta\theta^{*}}^{2}}{2\omega_{k}\omega_{k+1}}+\alpha\left(1-\frac{\tan\theta_{k+1}}{\tan\theta_{k}}\right)=0$ (37)
@@
$\mathcal{Z}\triangleq\begin{bmatrix}\theta_{k+1}-\theta_{k}-(-1)^{k+1}{\Delta\theta^{*}}\\ \dfrac{g{\Delta\theta^{*}}^{2}}{2\omega_{k}\omega_{k+1}}+\alpha\left(1-\dfrac{\tan\theta_{k+1}}{\tan\theta_{k}}\right)\end{bmatrix}$ (38)
@@ IV-B Stable Juggling
$\theta^{\mathrm{even}}=\pi-\theta^{\mathrm{odd}}$ (39)
i.e., the orientations at which impulsive inputs are applied on the devil-stick are symmetric about the vertical axis.
@@
$\frac{g{\Delta\theta^{*}}^{2}}{2\omega_{k}\omega_{k+1}}=\alpha\left(\frac{\tan\theta_{k+1}}{\tan\theta_{k}}-1\right)$ (40)
@@
$\frac{g{\Delta\theta^{*}}^{2}}{2\omega_{k+1}\omega_{k+2}}=\alpha\left(\frac{\tan\theta_{k}}{\tan\theta_{k+1}}-1\right)$ (41)
where (7) has been used to replace $\theta_{k+2}$ by $\theta_{k}$. Dividing (40) by (41) and simplifying, we obtain
$\omega_{k+2}=-\frac{\tan\theta_{k+1}}{\tan\theta_{k}}\omega_{k}$ (42)
@@
$\omega_{k+2}=\omega_{k}\quad\iff\quad\tan\theta_{k+1}=-\tan\theta_{k}\,\,\,\forall k$
To show sufficiency, observe that (5) and (39) imply
$\theta_{k+1}=\pi-\theta_{k}\ \forall k\quad\Rightarrow\quad\tan\theta_{k+1}=-\tan\theta_{k}\ \forall k$
To show necessity, observe that $\tan\theta_{k+1}=-\tan\theta_{k}$ can be solved as
$\theta_{k+1}=n\pi-\arctan(\tan\theta_{k}),\quad n\in\mathbb{Z}$
If $\theta_{k}=\theta^{\mathrm{odd}}$, $\arctan(\tan\theta_{k})=\theta^{\mathrm{odd}}$ since $\theta^{\mathrm{odd}}\in(0,\pi/2)$. It follows from (5) that $\theta_{k+1}=\theta^{\mathrm{even}}$, and $n=1$ provides the only feasible solution $\theta^{\mathrm{even}}=\pi-\theta^{\mathrm{odd}}$ since $\theta^{\mathrm{even}}\in(\pi/2,\pi)$. If $\theta_{k}=\theta^{\mathrm{even}}$, $\arctan(\tan\theta_{k})=\theta^{\mathrm{even}}-\pi$ since $\theta^{\mathrm{even}}\in(\pi/2,\pi)$. It follows from (5) that $\theta_{k+1}=\theta^{\mathrm{odd}}$, and $n=0$ provides the only feasible solution $\theta^{\mathrm{odd}}=\pi-\theta^{\mathrm{even}}$ since $\theta^{\mathrm{odd}}\in(0,\pi/2)$, which implies $\theta^{\mathrm{even}}=\pi-\theta^{\mathrm{odd}}$. ∎
@@
For the DVHC in (12), it is not possible to achieve stable juggling between orientations that are asymmetric about the vertical axis, i.e., $\theta^{\mathrm{even}}\neq\pi-\theta^{\mathrm{odd}}$.
Corollary 1 follows directly from (42). Theorem 1 implies that a unique stable, 2-periodic juggling motion corresponds to unique $\omega_{k}$ values for each $\theta_{k}\in\{\theta^{\mathrm{odd}},\theta^{\mathrm{even}}\}$, where $\theta^{\mathrm{odd}}$ and $\theta^{\mathrm{even}}$ satisfy (39); these values are:
$\omega_{k}=\begin{cases}\omega^{*},\qquad\quad\ \ k\,\ \mathrm{odd}\\ -\dfrac{g{\Delta\theta^{*}}^{2}}{4\omega^{*}\alpha},\quad k\,\ \mathrm{even}\end{cases}$ (43)
where $\omega^{*}<0$ may be chosen arbitrarily, and the value of $\omega_{k}$ corresponding to even $k$ is found from (37) with $\left(1-{\tan\theta_{k+1}}/{\tan\theta_{k}}\right)=2$, which follows from (39). Further, it follows from (30) and (43) that there exist the two unique times-of-flight
$\delta_{k}=\begin{cases}-{4\omega^{*}\alpha}/{(g{\Delta\theta^{*}})},\quad k\,\ \mathrm{odd}\\ -{{\Delta\theta^{*}}}/{\omega^{*}},\qquad\quad\ \,k\,\ \mathrm{even}\end{cases}$ (44)
From (31), (37), (39), and (43), we obtain
$I_{k}=\frac{(-1)^{k}2m\alpha}{{\Delta\theta^{*}}\cos\theta^{\mathrm{odd}}}\left[\omega^{*}+\frac{g{\Delta\theta^{*}}^{2}}{4\omega^{*}\alpha}\right]$ (45)
which indicates that the magnitude of impulse applied at either symmetric orientation of the devil-stick is the same, with opposite signs. From (32) and (39), we obtain
$r_{k}=\frac{J{\Delta\theta^{*}}\cos\theta^{\mathrm{odd}}}{2m\alpha}$ (46)
which implies that the location of application of the impulse is identical for either symmetric orientation of the devil-stick. These results demonstrate that the DZD (38) permits infinitely many stable, 2-periodic orbits when (39) holds, with each orbit corresponding to a unique choice of $\omega^{*}<0$.
IV-C Planar Symmetric Juggling
A special juggling motion, which is symmetric in both orientation and angular velocity about the vertical [8, 9] occurs when (39) holds and
$\omega^{*}=-\frac{{\Delta\theta^{*}}}{2}\sqrt{\frac{g}{\alpha}}\quad\Rightarrow\quad\omega_{k}=\begin{cases}\ \ \omega^{*},\quad k\,\ \mathrm{odd}\\ -\omega^{*},\quad k\,\ \mathrm{even}\end{cases}$ (47)
for which $\delta_{k}=-{\Delta\theta^{*}}/\omega^{*}$ $\forall k$. The steady-state solutions for states and control inputs in [8] are reproduced when $\omega^{*}$ is chosen as in (47) for the DVHC in (12).

### v2

@@ I Introduction
We consider the problem of planar juggling of a devil-stick, which is a nonprehensile manipulation task using impulsive inputs. The dynamics and control of juggling has been considered in [zavala-rio_control_1999, brogliato_control_2000, spong_impact_2001, buehler_planning_1994, sepulchre_stabilization_2003, sanfelice_hybrid_2007]. The combined control problem of the object and robot, with the associated unilateral constraints and impact laws, was considered in [zavala-rio_control_1999] for one-DOF ball juggling, and [brogliato_control_2000] for a class of complementary slackness jugglers. Other strategies for juggling point objects have appeared in [spong_impact_2001, buehler_planning_1994, sepulchre_stabilization_2003, sanfelice_hybrid_2007]. In contrast to point objects, juggling a devil-stick [schaal_open_1993, kant_non-prehensile_2021, kant_juggling_2022, khandelwal_nonprehensile_2023] requires modeling orientation in addition to position. The dynamics and closed-loop control design for planar symmetric juggling appeared in [kant_non-prehensile_2021, kant_juggling_2022]. A coordinate transformation was used to convert the orbital stabilization problem to one of fixed-point stabilization using a Poincaré map in a rotating reference frame. Consequently, the dynamic model only permitted orbits with states that are symmetric about the vertical axis.
We use discrete virtual holonomic constraints (DVHCs), introduced in [khandelwal_discrete_2025], to define the trajectory of the devil-stick at discrete instants in the inertial frame. Similar to VHCs [maggiore_virtual_2013, mohammadi_dynamic_2018, shiriaev_constructive_2005, grizzle_asymptotically_2001], DVHCs are geometric constraints on the generalized coordinates that avoid the need for tracking a time-varying reference trajectory, making the approach robust to unmodeled losses. However, while VHCs are enforced using continuous feedback such that they are satisfied for all time, DVHCs are enforced using discrete feedback and are satisfied only at the instants when impulsive inputs are applied. The DVHC approach applies to a wide class of underactuated mechanical systems with actuation capabilities restricted to impulsive inputs. This paper extends the concept of DVHCs from rotations to oscillations of the passive coordinate for the dynamical system in [khandelwal_discrete_2025]. Like the zero dynamics induced by VHCs, the discrete zero dynamics (DZD) induced by DVHCs can be used to infer nonlinear system stability.
The main contributions of this work, in comparison to [kant_non-prehensile_2021], are highlighted below.
@@
  • The DZD shows that the system permits infinitely many symmetric or asymmetric 2-periodic juggling motions, unlike [kant_non-prehensile_2021, kant_juggling_2022], which required juggling motions to be symmetric in both position and velocity about the vertical axis.
The impulse controlled Poincaré map (ICPM) approach [kant_orbital_2020, kant_juggling_2022] is adapted to stabilize any 2-periodic juggling motion.
This paper is organized as follows. In Section II, we define the planar devil-stick juggling problem and present the hybrid dynamics of the system. In Section III, we present the DVHC approach to designing the trajectory of the devil-stick, and a control design that enforces the DVHCs. In Section IV, we derive the DZD, which provides conditions for stable, periodic juggling. In Section V, we describe a method to stabilize a desired juggling motion. In Section VI, we validate the approach using simulations. The Appendix presents results for the special case when the devil-stick is juggled between orientations that are symmetric about the vertical axis.
@@ II System Dynamics
II-A System Description
Consider the three-DOF devil-stick in the vertical $xy$ plane in Fig. 1. The devil-stick has length $\ell$, mass $m$, and mass moment of inertia $J$ about its center-of-mass. Its configuration is described by the independent generalized coordinates $(h_{x},h_{y},\theta)$, where $h_{x},h_{y}$ denote the Cartesian coordinates of its center-of-mass, and $\theta$ denotes its orientation measured positive CCW with respect to the horizontal axis. It is juggled using the control inputs $I$ and $r$, where $I$ is the impulse applied normal to the devil-stick and $r$ is the distance of its point of application from the center-of-mass. The devil-stick is juggled between the orientations $\theta^{\mathrm{odd}}\in(0,\pi/2)$ and $\theta^{\mathrm{even}}\in(\pi/2,\pi)$, which are chosen a priori. Therefore, $\theta\in[\theta^{\mathrm{odd}},\theta^{\mathrm{even}}]$ $\forall t$ during the juggling task. Impulsive inputs are applied alternately at $\theta=\theta^{\mathrm{odd}}$ and $\theta=\theta^{\mathrm{even}}$, and the devil-stick falls freely under gravity between successive applications of impulsive inputs, making the resulting system dynamics hybrid. Since there is one fewer control input than the number of generalized coordinates, the system is underactuated with one passive DOF. The vector of generalized coordinates is
@@ II-A System Description
where $v\in\mathbb{R}^{2}$ and $\omega\in\mathbb{R}$; $v_{x},v_{y}$ denote the Cartesian components of the velocity of the center-of-mass of the devil-stick, and $\omega$ its angular velocity.
@@ II-B Hybrid Dynamic Model
The hybrid dynamic model captures a single impulsive actuation [^1] Impulsive inputs cause no change in position coordinates and finite jumps in velocities [kant_orbital_2020]., followed by a flight phase under gravity. The dynamic model in [khandelwal_discrete_2025], which applied to rotations of $\theta$, is adapted here for oscillations of $\theta$. Using $k$ to represent instants immediately prior to applications of impulsive inputs, the dynamics of the active positions and velocities are [khandelwal_discrete_2025]
@@
The dynamics of the passive states are [khandelwal_discrete_2025]
@@
Remark 1
Here, we do not model the robot manipulator; its motion must be controlled to realize the desired control inputs, ensuring that $\begin{bmatrix}\theta-\theta^{\mathrm{odd}}&\theta^{\mathrm{even}}-\theta\end{bmatrix}^{T}\geq 0$ $\forall t$ by considering the unilateral constraints between the devil-stick and robot - see the recursive strategy in [zavala-rio_control_1999, brogliato_control_2000].
@@ III-A Constraints on Positions
We use discrete virtual holonomic constraints (DVHCs) [khandelwal_discrete_2025] to define the trajectory of the center-of-mass of the devil-stick in terms of its orientation at the instants $k$ that impulsive inputs are applied on the devil-stick:
@@
$\Phi(\theta_{k})=\begin{bmatrix}\Phi_{x}(\theta_{k})&\Phi_{y}(\theta_{k})\end{bmatrix}^{T}$ (12)
where the DVHC components $\Phi_{x}$, $\Phi_{y}$ must be well-defined and nonsingular for each $\theta_{k}\in\{\theta^{\mathrm{odd}},\theta^{\mathrm{even}}\}$, and $\Phi_{x}$ must satisfy $\Phi_{x}(\theta^{\mathrm{odd}})>0$, $\Phi_{x}(\theta^{\mathrm{even}})<0$.
@@ III-C Control Design Enforcing the DVHC
We find the control inputs that enforce the DVHC following the approach in [khandelwal_discrete_2025]. Using (11) and (19) in (3), we obtain
@@
We define
$\eta_{k}\triangleq\Phi(\theta_{k+1})-\Phi(\theta_{k}),\quad\eta_{k}=\begin{bmatrix}\eta_{xk}&\eta_{yk}\end{bmatrix}^{T}$ (23)
and rewrite (21) as
$\begin{split}(\lambda-\mathbb{I}_{2})\rho_{k}+\eta_{k}=&[D\rho_{k}+\Psi(\theta_{k},\omega_{k})]\delta_{k}\\ &+\begin{bmatrix}-\sin\theta_{k}\\ \cos\theta_{k}\end{bmatrix}\frac{I_{k}\delta_{k}}{m}+\begin{bmatrix}0\\ -\frac{1}{2}g\delta_{k}^{2}\end{bmatrix}\end{split}$ (24)
The $x$ and $y$ components of (24) can be expressed separately as
@@
where $\rho_{k}=\begin{bmatrix}\rho_{xk}&\rho_{yk}\end{bmatrix}^{T}$, $D\rho_{k}=\begin{bmatrix}D\rho_{xk}&D\rho_{yk}\end{bmatrix}^{T}$, and $\Psi=\begin{bmatrix}\Psi_{x}&\Psi_{y}\end{bmatrix}^{T}$. Since $\sin\theta_{k}\neq 0$ for any $\theta_{k}$ in (5), we eliminate $I_{k}$ between (25a) and (25b) to obtain
@@
The time-of-flight $\delta_{k}$ is obtained by solving the above quadratic equation for the feasible value of $\delta_{k}$. The value of $I_{k}$ is then obtained from (25a) as
@@
It can be shown that [khandelwal_discrete_2025]
$D\rho_{k+1}=\frac{(\lambda-\mathbb{I}_{2})\rho_{k}}{\delta_{k}}$ (29)
@@
A dead-beat controller arises when $\lambda_{x}=\lambda_{y}=0$; however, this may not be feasible for the robot [brogliato_control_2000].
@@
$\delta_{k}$ $=\frac{(-1)^{k+1}2\omega_{k}(\eta_{xk}\cot\theta_{k}+\eta_{yk})}{g{\Delta\theta^{*}}}$ (30)
$I_{k}$ $=\frac{(-1)^{k+1}m\eta_{xk}}{\sin\theta_{k}}\left[\frac{\omega_{k}}{{\Delta\theta^{*}}}-\frac{g{\Delta\theta^{*}}}{2\omega_{k}(\eta_{xk}\cot\theta_{k}+\eta_{yk})}\right]$ (31)
$r_{k}$ $=-\frac{(-1)^{k+1}J{\Delta\theta^{*}}\sin\theta_{k}}{m\eta_{xk}}$ (32)
@@ IV-A Derivation
The discrete zero dynamics (DZD) provides the dynamics of passive coordinate $\theta$ and passive velocity $\omega$ when the DVHCs are satisfied, and system trajectories evolve such that $\rho_{k}\equiv 0\Rightarrow h(k)\equiv\Phi(\theta_{k})$ and $D\rho_{k}\equiv 0\Rightarrow v(k)\equiv\Psi(\theta_{k},\omega_{k})$. It follows from (19), (20), and (23) that
$v_{x}(k)$ $=-\frac{(-1)^{k}\omega_{k}\eta_{xk}}{{\Delta\theta^{*}}}$ (33a)
$v_{y}(k)$ $=-\frac{(-1)^{k}\omega_{k}\eta_{yk}}{{\Delta\theta^{*}}}-\dfrac{(-1)^{k}g{\Delta\theta^{*}}}{2\omega_{k}}$ (33b)
Similarly, it follows from (7), (19), (20), and (23) that
$v_{x}(k+1)$ $=-\frac{(-1)^{k}\omega_{k+1}\eta_{xk}}{{\Delta\theta^{*}}}$ (34a)
$v_{y}(k+1)$ $=-\frac{(-1)^{k}\omega_{k+1}\eta_{yk}}{{\Delta\theta^{*}}}+\dfrac{(-1)^{k}g{\Delta\theta^{*}}}{2\omega_{k+1}}$ (34b)
since $\eta_{k+1}=\Phi(\theta_{k+2})-\Phi(\theta_{k+1})=-[\Phi(\theta_{k+1})-\Phi(\theta_{k})]=-\eta_{k}$. The input $I_{k}$ can be eliminated between the two equations in (4) to obtain
@@
$\eta_{xk}\cot\theta_{k}+\eta_{yk}-\frac{g{\Delta\theta^{*}}^{2}}{2\omega_{k}\omega_{k+1}}=0$ (37)
@@
$\mathcal{Z}\triangleq\begin{bmatrix}\theta_{k+1}-\theta_{k}-(-1)^{k+1}{\Delta\theta^{*}}\\ \eta_{xk}\cot\theta_{k}+\eta_{yk}-\dfrac{g{\Delta\theta^{*}}^{2}}{2\omega_{k}\omega_{k+1}}\end{bmatrix}$ (38)
@@ IV-B Stable Juggling
$\Phi_{y}(\theta_{k})=\beta-\sigma\Phi_{x}(\theta_{k}),\quad\sigma\triangleq\frac{1}{2}(\cot\theta^{\mathrm{odd}}+\cot\theta^{\mathrm{even}})$ (39)
where $\sigma$ is a constant, and $\beta\in\mathbb{R}$ is an arbitrary constant.
@@
$\frac{g{\Delta\theta^{*}}^{2}}{2\omega_{k}\omega_{k+1}}=\eta_{xk}\cot\theta_{k}+\eta_{yk}$ (40)
@@
$\frac{g{\Delta\theta^{*}}^{2}}{2\omega_{k+1}\omega_{k+2}}=-\eta_{xk}\cot\theta_{k+1}-\eta_{yk}$ (41)
since $\eta_{k+1}=-\eta_{k}$ - see the text after (34). Dividing (40) by (41) and simplifying, we obtain
$\omega_{k+2}={-\left(\frac{\eta_{xk}\cot\theta_{k}+\eta_{yk}}{\eta_{xk}\cot\theta_{k+1}+\eta_{yk}}\right)}\omega_{k}$ (42)
@@
$\omega_{k+2}=\omega_{k}$
$\iff$ $\eta_{xk}\cot\theta_{k}+\eta_{yk}=-\eta_{xk}\cot\theta_{k+1}-\eta_{yk}$
$\iff$ $\frac{1}{2}(\cot\theta_{k}+\cot\theta_{k+1})\eta_{xk}+\eta_{yk}=0$
Since $\theta_{k}=\theta^{\mathrm{odd}}\Rightarrow\theta_{k+1}=\theta^{\mathrm{even}}$, and $\theta_{k}=\theta^{\mathrm{even}}\Rightarrow\theta_{k+1}=\theta^{\mathrm{odd}}$, $\frac{1}{2}(\cot\theta_{k}+\cot\theta_{k+1})=\sigma\,\forall k$, with $\sigma$ defined in (39). Thus,
$\omega_{k+2}=\omega_{k}$
$\iff$ $\sigma\eta_{xk}+\eta_{yk}=0$
$\iff$ $\sigma\Phi_{x}(\theta_{k+1})+\Phi_{y}(\theta_{k+1})=\sigma\Phi_{x}(\theta_{k})+\Phi_{y}(\theta_{k})$
$\iff$ $\sigma\Phi_{x}(\theta_{k})+\Phi_{y}(\theta_{k})=\beta\quad\forall k$
where $\beta$ is a constant. This completes the proof. ∎
@@
If $\theta^{\mathrm{even}}=\pi-\theta^{\mathrm{odd}}$, i.e., the devil-stick is juggled between orientations symmetric about the vertical axis, the DVHC (12) satisfies $\Phi_{y}(\theta_{k})=\beta$.
Proof:
The proof follows directly from (39), since $\theta^{\mathrm{even}}=\pi-\theta^{\mathrm{odd}}\Rightarrow\sigma=0$. ∎
Theorem 1 implies that a unique stable, 2-periodic juggling motion corresponds to two unique $\omega_{k}$ values for each $\theta_{k}\in\{\theta^{\mathrm{odd}},\theta^{\mathrm{even}}\}$, when the DVHC satisfies (39). These values of $\omega_{k}$ are:
$\omega_{k}$ $=\begin{cases}\omega^{*},\quad\quad\ \ k\,\ \mathrm{odd}\\ \dfrac{g{\Delta\theta^{*}}^{2}}{2\omega^{*}\xi},\quad k\,\ \mathrm{even}\end{cases}$ (43)
$\xi$ $\triangleq-\frac{1}{2}(\cot\theta^{\mathrm{odd}}-\cot\theta^{\mathrm{even}})\eta_{x}^{*}$ (44)
$\eta_{x}^{*}$ $\triangleq\Phi_{x}(\theta^{\mathrm{odd}})-\Phi_{x}(\theta^{\mathrm{even}})$ (45)
where $\xi<0$, $\eta_{x}^{*}>0$ are constants, and $\omega^{*}<0$ may be chosen arbitrarily; the value of $\omega_{k}$ corresponding to even $k$ is found from (37) and (39). Further, it follows from (30) and (43) that there exist the two unique times-of-flight
$\delta_{k}=\begin{cases}{2\omega^{*}\xi}/{(g{\Delta\theta^{*}})},\quad k\,\ \mathrm{odd}\\ -{{\Delta\theta^{*}}}/{\omega^{*}},\qquad\ \,k\,\ \mathrm{even}\end{cases}$ (46)
From (31), (39), and (43), we obtain
$I_{k}=\frac{(-1)^{k}m\eta_{x}^{*}}{\sin\theta_{k}}\left[\frac{\omega^{*}}{{\Delta\theta^{*}}}-\frac{g{\Delta\theta^{*}}}{2\omega^{*}\xi}\right]$ (47)
From (32) and (39), we obtain
$r_{k}=\frac{J{\Delta\theta^{*}}\sin\theta_{k}}{m\eta_{x}^{*}}$ (48)
These results demonstrate that the DZD (38) permits infinitely many stable, 2-periodic orbits between arbitrary $\theta^{\mathrm{odd}}\in(0,\pi/2)$, $\theta^{\mathrm{even}}\in(\pi/2,\pi)$ when the DVHC (12) is chosen such that (39) holds, with each orbit corresponding to a unique choice of $\omega^{*}<0$.
Remark 5
Identical trajectories of the center-of-mass between successive impacts are obtained when $\delta_{k}$ is constant $\forall k$ in (46), i.e., ${2\omega^{*}\xi}/{(g{\Delta\theta^{*}})}=-{{\Delta\theta^{*}}}/{\omega^{*}}$, which implies
$\omega^{*}=-{\Delta\theta^{*}}\sqrt{-\frac{g}{2\xi}}\ \Rightarrow\ \omega_{k}=\begin{cases}\ \ \omega^{*},\quad k\,\ \mathrm{odd}\\ -\omega^{*},\quad k\,\ \mathrm{even}\end{cases}$ (49)